Angle A and angle B are supplementary angles. If the measure of angle B is 6 degrees greater twice the measure of angle A, find the measure of each angle.
Answer by nerdybill(7384)

Angle A and angle B are supplementary angles. If the measure of angle B is 6 degrees greater twice the measure of angle A, find the measure of each angle.
.
By definition of "supplementary angles" we have:
A + B = 180
.
Then, from:
"angle B is 6 degrees greater twice the measure of angle A" we have
B = 2A+6
.
Substitute the above into:
A + B = 180
A + 2A+6 = 180
3A+6 = 180
3A = 174
A = 58 degrees
.
Substitute above into the following to find B:
A + B = 180
58 + B = 180
B = 122 degrees